2. Convert to i) cylindrical and ii) Spherical coordinates. 319-x2 /9-x2 - y2 V x2 + y2 + z2 dz dy dx

convert the integrals from poolar coordinates to those 2 coordinates in the question!
Answer #1

9-x2 y2 9-02 I = hmit )2가y주229 y=J9-x to Z=0 In Spheucol Coordingtes put Sinf Comp zEf Corp 1 POpdppurd =zphpep Po B-O PE Cyl